OpenStudy (anonymous):

Find the ratio of x to y. x/5=2/3=5/y

OpenStudy (freckles):

well let's solve our two equations first: \[\frac{x}{5}=\frac{2}{3} \text{ and } \frac{2}{3}=\frac{5}{y}\]

OpenStudy (freckles):

can you solve the first equation for x?

OpenStudy (anonymous):

x=0.3.??

OpenStudy (xapproachesinfinity):

leave it as 10/3 and that won't be 0.3 anyway \[x=\frac{10}{3} \\ y=\frac{15}{2}\] now take the ratio \[\frac{x}{y}=?\]